OpenStudy (yoloshroom):

Solve exactly for X... Log(x+3)-log(x+2)=log(4)

OpenStudy (yoloshroom):

well first you do, log((x+3)/(x+2))=log(4) but then what?

OpenStudy (anonymous):

raise each side to the power of 10

OpenStudy (anonymous):

or equate the arguments

OpenStudy (mathstudent55):

Use the rules of logs: \(\log a + \log b = \log ab\) \(\log_b x = y \iff b^y = x\) If \(\log a = \log b\) then \(a = b\)

OpenStudy (mathstudent55):

\(\log a - \log b = \log \dfrac{a}{b} \)

OpenStudy (mathstudent55):

Correct.

OpenStudy (yoloshroom):

ahhhhh good. i thought so, @mathstudent55 I just wasnt sure..

OpenStudy (mathstudent55):

Look at the third rule above.

OpenStudy (mathstudent55):

\(\log \dfrac{x+3}{x+2} = \log 4\) Now apply: \(\log a = \log b\), then \(a = b\)

OpenStudy (yoloshroom):

but my answer comes out to -5/3=x right? doesnt that not work because like -5/3 is less than the base?

OpenStudy (yoloshroom):

or am i just confusing myself?

OpenStudy (mathstudent55):

Once the logs are equal, then \(\dfrac{x+3}{x+2} = 4\) Did you solve this equation?

OpenStudy (yoloshroom):

-5/3 ^_^

OpenStudy (yoloshroom):

would that be my final answer?

OpenStudy (mathstudent55):

\(\dfrac{x+3}{x+2} = 4\) \(x + 3 = 4(x + 2)\) \(x + 3 = 4x + 8\) \(3x = -5\) \(x = - \dfrac{5}{3} \) Exactly. Good job!

OpenStudy (yoloshroom):

i got one more, you up for helping?

OpenStudy (mathstudent55):

Sure. Pls start a new post.